OpenStudy (anikate):

simplify: (x+y)^2 -1/4

OpenStudy (anonymous):

x^2+2xy+y^2-1/4

OpenStudy (anikate):

that's it?

OpenStudy (anonymous):

\[\left(x+y+\frac{1}{2}\right)\left(x+y-\frac{1}{2}\right) \]